# The Relational Origins of Rules in Online Communities

### v1

@@ Abstract.
Where do rules come from in online communities? While prior studies of online community governance in social computing have sought to characterize rules by their functions within communities and documented practices of rule enforcement, they have largely overlooked rule adoption and change. This study investigates how and why online communities adopt and change their rules. We conducted a grounded theory-based analysis of 40 in-depth interviews with community leaders from subreddits, Fandom wikis, and Fediverse servers, and identified seven processes involved in the adoption of online community rules. Our findings reveal that, beyond regulating behavior and solving functional intra-community problems, rules are also adopted and changed for relational reasons, such as signaling or reinforcing community legitimacy and identity to other communities. While rule change was often prompted by challenges during community growth or decline, change also depended on volunteer leaders’ work capacity, the presence of member feedback mechanisms, and relational dynamics between leaders and members. The findings extend prior theories from social computing and organizational research, illustrating how institutionalist and ecological explanations of the relational origins of rules complement more functional accounts. The results also support design recommendations that integrate the relational aspects of rules and rulemaking to facilitate successful governance across communities’ lifecycles.
copyright: ccjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Preprint under review; May, 2025; posted to https://arxiv.org
@@ 1. Introduction
Rules serve many purposes in online communities. Rules codify social norms (Chandrasekharan et al., 2018; Fiesler et al., 2018), signal appropriate behavior (Matias, 2019b), and are subjects of debate in online governance (Gillespie, 2018; Stockinger et al., 2023; Jhaver et al., 2019c; Crawford and Gillespie, 2016). Prior studies in social computing offer insights into extant rules (e.g., Fiesler et al., 2018; Nicholson et al., 2023; Hwang and Shaw, 2022; Keegan and Fiesler, 2017; Reddy and Chandrasekharan, 2023) as well as practices of rule enforcement (e.g., Dosono and Semaan, 2019; Wohn, 2019; Cullen and Kairam, 2022; Matias, 2019b), including the impacts of automated moderation and enforcement tools (e.g., Chandrasekharan et al., 2019; TeBlunthuis et al., 2021; Halfaker and Geiger, 2020). However, most studies of rules take for granted that rules must also be designed and adopted, yet we know little about the origins of rules in online communities. Understanding the origins of rules is important because it reveals a design space of rulemaking and community governance.
@@
There are two broad approaches to understanding rulemaking in the social computing literature. First, a top-down “social engineering” perspective (Kraut and Resnick, 2012) aims to identify and evaluate generalizable principles that platforms can implement to regulate behavior and guide effective content moderation (Chandrasekharan et al., 2019, 2022; Matias, 2019b; Jhaver et al., 2021; Seering et al., 2017; Horta Ribeiro et al., 2021). An alternative “institutional analysis and development” (IAD) or ”institutional design” perspective (Frey et al., 2019b) emphasizes the bottom-up and localized development of rules, rulemaking norms, and enforcement practices within communities (e.g., Schneider et al., 2021). Despite their divergent orientations, both approaches assume that rules are primarily functional—i.e., created to solve problems located in the community adopting a rule.
@@
Our alternative approach begins from the observation that many online communities share similar rules (Reddy and Chandrasekharan, 2023; Fiesler et al., 2018; Nicholson et al., 2023; Hwang and Shaw, 2022). A functional account of rulemaking would attribute these patterns to common problems communities face. However, research in organizational sociology has rejected such explanations of rule similarity in firms since the 1970s. This work—including “new institutionalism” (e.g., Meyer and Rowan, 1977; DiMaggio and Powell, 1983; Tolbert and Zucker, 1983) as well as “organizational ecology” (Hannan and Freeman, 1984; Carroll, 1985)—has argued that similarities in organizational structures are best explained in reference to external social pressures. We apply these insights to the crafting of the rules in online communities.
@@
To understand the origins of rules, we conducted 40 in-depth, open-ended interviews with leaders of communities from three different sociotechnical systems: subreddit communities on Reddit, wikis on Fandom, and servers in the Fediverse. Interviewees were recruited from clusters of communities with similar rules in each setting. We report findings based on a comparative, grounded theory analysis (Charmaz, 2015) that draws on the earlier work in organizational sociology as sensitizing theories.
@@
We find that rules originate and change for a mix of functional and relational purposes that go beyond merely providing the basis for enforcement decisions or conveying community goals and norms. Rules and rulemaking play important, underappreciated roles as communities navigate relationships with internal and external constituencies and constraints. In particular, we find that community leaders use rules to draw distinctions between communities, manage issues of community growth as well as decline, and legitimize the community to external peers and platforms.
@@
Our paper makes several contributions to social computing theory focused on governance. First, our account of the origins of rules emphasizes how rules mediate online communities’ relationships to other communities, platform rules and affordances, and audiences of potential and current members, all of which constrain rulemaking. This relational aspect of rulemaking extends prior functionalist accounts and expands understanding of the design space of digital institutions and community self-governance (Frey et al., 2019a; Zhang et al., 2020; Frey and Schneider, 2021; Jhaver et al., 2021; Schneider, 2022). Additionally, we synthesize concepts from new institutionalism and organizational ecology to show how rules play a relational role in community ecosystems by communicating niche identity and legitimacy. Our work also makes empirical contributions with practical importance by filling a critical qualitative gap on the origins and practices of rule creation and change in online communities and highlight opportunities for rule adoption strategies across community lifecycle phases. We also provide comparative empirical evidence across three different systems to show how technological design can influence rule adoption and change.
@@ 2. Background
Rules are central to governance in online communities. Yet, when research in social computing has considered rules, it typically takes for granted how rules are initially adopted and how they evolve. Our paper seeks to build new theory about why and how rules in online communities are created and changed. Although our research design is inductive and our findings emerge from our data, there are two lines of existing scholarship that frame our study. The first line includes two major paradigms in prior social computing work on online community governance: the “social engineering” and “institutional design” perspectives. This scholarship provides the rationale for our study as well as its design. The second line of research comes from the “new institutionalist” and “organizational ecology” traditions in sociology. These literatures help us interpret and synthesize the themes that emerged from our analysis.
@@ 2.1. Governance and Rules in Online Communities
Governance is the social practices and processes that shape behavior within and between institutions (Bednar, 2023). As part of governance, rules are the explicit “norms, regulations, and expectations that regulate the behavior of individuals and interactions among them.” (March et al., 2000, p. 5). Social computing research has explored rules and governance in online communities—the virtual spaces where individuals engage around shared interests or goals through socially regulated interaction on a social media platform (Jones, 1997; Preece, 2000; Leimeister et al., 2004). [^1] For our purposes, a “social media platform” is a website that affords the creation of communities around user-generated content and interactions (Carr and Hayes, 2015) This study responds to two influential perspectives in social computing on rules in online communities.
The first is the top-down “social engineering” approach that centers the role of “designers,” and is exemplified by the work in Building Successful Online Communities (Kraut and Resnick, 2012). This research paradigm seeks to identify generalizable principles by which platform designers can shape behavior in ways that correspond to their goals. For example, Kraut and Resnick (2012)’s chapter on norms and rules lists a series of “design principles” that offer specific guidance for how a community’s designer can set injunctive norms (i.e., rules) to shape participants’ behavior. Other studies test how rule placement can increase salience and compliance (Matias, 2019b), evaluate how post-enforcement rule explanations impact subsequent user behavior (Jhaver et al., 2019c), and investigate the effectiveness of different rule enforcement mechanisms (Horta Ribeiro et al., 2021; Chandrasekharan et al., 2017). A significant line of research also focuses on describing existing governance approaches and developing new ones (Geiger, 2014; Seering et al., 2017; Chandrasekharan et al., 2019; Jhaver et al., 2019a, b; Zhang et al., 2020).
@@
Existing work on online communities has often approached the topic of rule change from the perspective of contestation. Scholars note how moments of rule enforcement can become opportunities to contest the meaning and understanding of rules (de Laat, 2012; Crawford and Gillespie, 2016; Matei and Dobrescu, 2010; Beschastnikh et al., 2008; Kriplean et al., 2007). That said, the little empirical evidence we have on the subject suggests that written rules in online communities change infrequently. For example, Hwang and Shaw (2022) finds that rules on the five largest language editions of Wikipedia are increasingly less edited even as they continue to be discussed. Evidence from the broader social computing literature offers some explanations for why rule changes may not be frequent online. Online communities are numerous and relatively cheap to reproduce (Benkler, 2006). This might make opportunities for rule change less likely if members can simply leave and join another community that already has the rules that they want. Hirschman (1970) suggests that the presence of alternatives can influence whether leaders ignore problems or enact change. Moreover, scholars have noted that opportunities for “voice” in online communities are limited (Schneider, 2022; Frey and Schneider, 2021; Shaw and Hill, 2014).
A key assumption of both the social engineering and institutional design perspectives is that rules are functional and that rule selection and change can be understood as the output of rational processes: choices to adopt or change rules are determined by specific problems within specific communities. From a functional standpoint, efforts to understand governance in both perspectives focused on evaluating whether rules or rulemaking successfully mitigate internal problems. For example, research on content moderation (Grimmelmann, 2017) across the two perspectives is centrally concerned with the effectiveness (Chandrasekharan et al., 2019) and consequences (Jhaver et al., 2019a) of rule enforcement practices. Both perspectives treat rules as emerging to address these internal problems, just from different sources. To the extent that communities wind up with similar rules, it is because they have similar problems that benefit from similar solutions. The sources of rules have largely remained unexamined in both streams of research.
@@ 2.2. Sociological Perspectives on Organizational Similarity
The most important body of scholarship in organizational sociology we rely on is known as the new institutionalism. It was first articulated by Meyer and Rowan (1977) and DiMaggio and Powell (1983). Confusingly, there are several scholarly traditions in economics, political science, and sociology referred to as “institutionalism” and even “new institutionalism” (Powell and Dimaggio, 1991). These traditions have distinct emphases and often define institutions differently. For example, the institutional design perspective described in §2.1 draws from North (1990), Ostrom (1990), and others associated with what is sometimes called “new institutional economics.” This approach defines institutions as “rules of the game” and seeks to increase attention to the way written and unwritten rules shape and constrain the behavior of individuals and organizations.
@@
New institutionalism in sociology is a distinct and more conceptually narrow line of work that argues that organizations’ rules and routines do not reflect the rational solution to problems faced by organizations, something taken for granted in much of the work on rules in economics. They are instead based on “myths” (Meyer and Rowan, 1977) and are the results of isomorphic pressures that cause organizations to adopt similar structures, rules, and routines as a way to earn social legitimacy (DiMaggio and Powell, 1983). DiMaggio and Powell (1983) argue that organizations will copy routines from other successful groups (mimetic isomorphism), adopt them because they reflect expectations from groups like professions (normative isomorphism), or choose them because outsiders like investors or governments require them (coercive isomorphism). Although specific sources of isomorphic pressure differ across organizations and environments, this line of scholarship provides an obvious entry point to approach our observation of similarity in online community rules. After all, online communities on a given platform face similar constraints because they are on the same social computing system (Caplan and boyd, 2018), are subject to similar external policies (Tosch et al., 2024), and rely on similar tools (Hwang et al., 2024).
@@
A distinct tradition within organizational ecology treats organizational behavior as contingent on environmental factors that can promote either variation or conformity in organizational structure (Cyert and March, 1963; Hannan and Freeman, 1977). This “organizational ecology” perspective seeks to explain how organizations may differentiate themselves to avoid directly competing with one another and to meet diverse and specialized needs of audiences or markets (Carroll and Swaminathan, 2000). While the new institutionalists traditionally focused on explaining organizational similarity (i.e., why are organizations so similar?), ecologists more typically focused on variation (i.e., why are organizations so different?). The organizational ecology approach has been echoed by social computing scholars who note that communities can range immensely in their topics, scales, goals, and norms (Chandrasekharan et al., 2018; Weld et al., 2021; TeBlunthuis and Hill, 2022; Zhu et al., 2014, 2013; Hwang and Foote, 2021). Drawing explicitly from organizational ecology, TeBlunthuis et al. (2022) observes how normative differences between communities with similar topics enable communities to be mutualistic rather than competitive with one another. Related work has mapped out “ecologies” of the rules and values that would underpin such normative distinctions (Chandrasekharan et al., 2018; Weld et al., 2021; Fiesler et al., 2018; Nicholson et al., 2023; Hecht and Gergle, 2010). Although researchers have discussed deliberative processes and bureaucratization around rules within specific communities (Black et al., 2008; Butler et al., 2008), we do not know of work that seeks to explain how ecologies of community rules form in online contexts.
@@
An important distinction between the new institutionism and ecology traditions is in the way they approach change. For new institutionalists, organizational change is a central object of study. For example, a new law or regulation may pressure a firm to change its rules to comply with new legal requirements. Similarly, a nonprofit might change its rules to qualify for funding or to conform to published best practices. A key implication from the new institutionalist perspective is that organizations will face similar pressures that will make them structurally similar over time. Conversely, organizational ecologists rely heavily on the concept of structural inertia, which assumes that organizational change is difficult or even impossible (Hannan and Freeman, 1977; Young, 1988). If organizations appear similar, it is because they face similar selection pressures, which have caused dissimilar organizations to fail, or because there are positive returns to legitimacy that stem from similar organizations in an environment. That said, some organizational ecologists have argued that certain features of environments may lead to more organizational change (Hannan and Freeman, 1984).
@@
Although the new institutionalist and organizational ecology perspectives are quite different in their assumptions and predictions, both perspectives share a key sociological insight that is largely absent from most scholarship on governance in social computing: the origins of organizational forms—including rules—lie not only in function but also in the relationships between the organizations and their environments. According to both perspectives, we cannot fully understand communities by looking within communities or by focusing on the problems that the text of the rules purports to solve. In both perspectives, rules are deeply social and communicative.

### v2

@@ Abstract.
Where do rules come from in online communities? While prior studies of online community governance in social computing have sought to characterize rules by their functions and enforcement practices, scholars have largely overlooked rule adoption and change. This study investigates how and why online communities adopt and change their rules. We conducted a grounded theory-based analysis of 40 in-depth interviews with community leaders from subreddits, Fandom wikis, and Fediverse servers, and identified seven processes involved in the adoption of online community rules. Our findings reveal that, beyond functional reasons like regulating behavior and solving problems, rules are also adopted and changed for relational reasons, such as signaling or reinforcing community legitimacy and identity to other communities. While rule change was often prompted by challenges during community growth or decline, change also depended on volunteer leaders’ work capacity, the presence of member feedback mechanisms, and relational dynamics between leaders and members. The findings extend prior theories from social computing and organizational research, illustrating how institutionalist and ecological explanations of the relational origins of rules complement functional accounts. The results also support design recommendations that integrate the relational aspects of rules and rulemaking to facilitate successful governance across communities’ lifecycles.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Rules serve many purposes in online communities. Rules codify social norms (Chandrasekharan et al., 2018; Fiesler et al., 2018), signal appropriate behavior (Matias, 2019b), and are the subject of debate (Gillespie, 2018; Stockinger et al., 2023; Jhaver et al., 2019c; Crawford and Gillespie, 2016). Prior studies in social computing offer insights into extant rules (e.g., Fiesler et al., 2018; Nicholson et al., 2023; Hwang and Shaw, 2022; Keegan and Fiesler, 2017; Reddy and Chandrasekharan, 2023) as well as practices of rule enforcement (e.g., Dosono and Semaan, 2019; Wohn, 2019; Cullen and Kairam, 2022; Matias, 2019b), including the impacts of automated moderation and enforcement tools (e.g., Chandrasekharan et al., 2019; TeBlunthuis et al., 2021; Halfaker and Geiger, 2020). However, because most studies of rules take for granted that rules must also be designed and adopted, we know little about the origins of rules in online communities. How and why do rules come to be in online communities? What challenges do leaders face in creating and changing them?
@@
There are two broad approaches to understanding rulemaking in the social computing literature. First, a top-down “social engineering” perspective (Kraut et al., 2012) aims to identify and evaluate generalizable principles that platforms can implement to regulate behavior and guide effective content moderation (Chandrasekharan et al., 2019, 2022; Matias, 2019b; Jhaver et al., 2021; Seering et al., 2017; Horta Ribeiro et al., 2021). An alternative “institutional analysis and development” (IAD) or ”institutional design” perspective (Frey et al., 2019b) emphasizes the bottom-up and localized development of rules, rulemaking norms, and enforcement practices within communities (e.g., Schneider et al., 2021). Despite their differences, both approaches assume that rules are primarily functional—i.e., created to solve problems in the community adopting a rule.
@@
Our alternative approach begins from the observation that many online communities share similar rules (Reddy and Chandrasekharan, 2023; Fiesler et al., 2018; Nicholson et al., 2023; Hwang and Shaw, 2022). A functional account of rulemaking would attribute these patterns to common problems communities face. However, studies of firms in organizational sociology has rejected such explanations. This work—including “new institutionalism” (e.g., Meyer and Rowan, 1977; DiMaggio and Powell, 1983; Tolbert and Zucker, 1983) as well as “organizational ecology” (Hannan and Freeman, 1984; Carroll, 1985)—has argued that similarities in organizational structures are best explained by external social pressures. We apply these insights to the crafting of the rules in online communities.
@@
To understand the origins of rules, we conducted 40 in-depth, open-ended interviews with leaders of communities from three different sociotechnical systems: subreddit communities on Reddit, wikis on Fandom, and servers in the Fediverse. Interviewees were recruited from clusters of communities with similar rules in each setting. We report findings based on a comparative, grounded theory analysis (Charmaz, 2015) that treats earlier work in organizational sociology as sensitizing theories.
@@
We find that rules originate and change for a mix of functional and relational purposes that go beyond merely providing the basis for enforcement decisions or conveying community goals and norms. We show that rules and rulemaking play important, underappreciated roles as communities navigate relationships with internal and external constituencies and constraints. In particular, we find that community leaders use rules to draw distinctions between communities, manage issues of community growth as well as decline, and legitimize their communities.
@@
Our paper makes several contributions to social computing theory focused on governance. First, our account of the origins of rules emphasizes how rules mediate online communities’ relationships to other communities, platform rules and affordances, and audiences of potential and current members, all of which constrain rulemaking. This relational aspect of rulemaking extends prior functionalist accounts and expands understanding of the design space of digital institutions and community self-governance (Frey et al., 2019a; Zhang et al., 2020; Frey and Schneider, 2021; Jhaver et al., 2021; Schneider, 2022). Additionally, we synthesize concepts from new institutionalism and organizational ecology to show how rules play a relational role in community ecosystems by communicating niche identity and legitimacy. We make empirical contributions by filling a critical qualitative gap on the origins and practices of rule creation and change in online communities with comparative evidence across three different systems. Finally, our findings highlight opportunities for rule adoption strategies across community lifecycle phases and show how technological design can influence rule adoption and change.
@@ 2. Background
Rules are central to governance in online communities. Yet, past research takes for granted how rules are initially adopted and how they evolve. Our paper seeks to build new theory about why and how rules in online communities are created and changed. Although our research design is inductive and our findings emerge from our data, we use two lines of existing scholarship to frame our study. The first line includes two major paradigms in prior social computing work on online community governance: the “social engineering” and “institutional design” perspectives. This scholarship provides the rationale for our study as well as its design. The second line of research comes from the “new institutionalist” and “organizational ecology” traditions in sociology. These literatures help us interpret and synthesize the themes that emerged from our analysis.
@@ 2.1. Governance and Rules in Online Communities
Governance is the social practices and processes that shape behavior within and between institutions (Bednar, 2023). As part of governance, rules are the explicit “norms, regulations, and expectations that regulate the behavior of individuals and interactions among them.” (March et al., 2000, p. 5). Social computing research has explored rules and governance in online communities—the virtual spaces where individuals engage around shared interests or goals through socially regulated interaction on a social media platform (Jones, 1997; Preece, 2000; Leimeister et al., 2004). [^1] For our purposes, a “social media platform” is a website that affords the creation of communities around user-generated content and interactions (Carr and Hayes, 2015). This study responds to two influential perspectives in social computing on rules in online communities.
The first is the top-down “social engineering” approach that centers the role of “designers,” and is exemplified by the work in Kraut and Resnick’s Building Successful Online Communities (Kraut et al., 2012). This research paradigm seeks to identify generalizable principles by which platform designers can shape behavior in ways that correspond to their goals. For example, Kraut et al. (2012)’s chapter on norms and rules lists a series of “design principles” that offer specific guidance for how a community’s designer can use rules to shape participants’ behavior. Other studies test how rule placement can increase salience and compliance (Matias, 2019b), evaluate how post-enforcement rule explanations impact subsequent user behavior (Jhaver et al., 2019c), and investigate the effectiveness of different rule enforcement mechanisms (Horta Ribeiro et al., 2021; Chandrasekharan et al., 2017). A significant line of research also focuses on describing existing governance approaches and developing new ones (Geiger, 2014; Seering et al., 2017; Chandrasekharan et al., 2019; Jhaver et al., 2019a, b; Zhang et al., 2020).
@@
Existing work on online communities has often approached the topic of rule change from the perspective of contestation. Scholars note how moments of rule enforcement can become opportunities to contest the meaning and understanding of rules (de Laat, 2012; Crawford and Gillespie, 2016; Matei and Dobrescu, 2010; Beschastnikh et al., 2008; Kriplean et al., 2007). That said, the little empirical evidence we have on rule change suggests that written rules in online communities change infrequently. For example, Hwang and Shaw (2022) finds that rules on the five largest language editions of Wikipedia are decreasingly edited even as they continue to be discussed. Evidence from the broader social computing literature offers some explanations for why rule changes may not be frequent. Online communities are numerous and relatively cheap to reproduce (Benkler, 2006). Hirschman (1970) suggests that the presence of alternatives can influence whether leaders ignore problems or enact change. Moreover, scholars have noted that opportunities for “voice” in online communities are limited (Schneider, 2022; Frey and Schneider, 2021; Shaw and Hill, 2014).
A key assumption of both the social engineering and institutional design perspectives is that rules are functional and that rule selection and change can be understood as the output of rational processes: choices to adopt or change rules are determined by specific problems within specific communities. From a functional standpoint, efforts to understand governance in both perspectives have focused on evaluating whether rules or rulemaking successfully mitigate internal problems. For example, research on content moderation (Grimmelmann, 2017) across the two perspectives is centrally concerned with the effectiveness (Chandrasekharan et al., 2019) and consequences (Jhaver et al., 2019a) of rule enforcement practices. The sources of rules have largely remained unexamined in both streams of research.
@@ 2.2. Sociological Perspectives on Organizational Similarity
The most important body of scholarship in organizational sociology we rely on is known as the new institutionalism. It was first articulated by Meyer and Rowan (1977) and DiMaggio and Powell (1983). Confusingly, there are several scholarly traditions in economics, political science, and sociology referred to as “institutionalism” and even “new institutionalism” (Powell and Dimaggio, 1991). These traditions have distinct emphases and often define institutions differently. For example, the institutional design perspective described in §2.1 draws from North (1990), Ostrom (1990), and others associated with what is sometimes called “new institutional economics.” This approach defines institutions broadly as “rules of the game” and seeks to increase attention to the way written and unwritten rules shape and constrain the behavior of individuals and organizations.
@@
New institutionalism in sociology is a distinct and more conceptually narrow line of work that argues that organizations’ rules and routines do not reflect the rational solution to problems faced by organizations—something taken for granted in much of the work on rules in economics. They are instead based on “myths” (Meyer and Rowan, 1977) and are the results of isomorphic pressures that cause organizations to adopt similar structures, rules, and routines as a way to earn social legitimacy (DiMaggio and Powell, 1983). DiMaggio and Powell (1983) argue that organizations will copy routines from other successful groups (mimetic isomorphism), adopt them because they reflect expectations from groups like professions (normative isomorphism), or choose them because outsiders like investors or governments require them (coercive isomorphism). Although specific sources of isomorphic pressure differ across organizations and environments, this line of scholarship provides an obvious entry point to approach our observation of similarity in online community rules. After all, online communities on a given platform face similar constraints because they are on the same social computing system (Caplan and boyd, 2018), are subject to similar external policies (Tosch et al., 2024), and rely on similar technology (Hwang et al., 2024).
@@
A distinct tradition within organizational ecology treats organizational behavior as contingent on environmental factors that can promote either variation or conformity in organizational structure (Cyert and March, 1963; Hannan and Freeman, 1977). This “organizational ecology” perspective seeks to explain how organizations may differentiate themselves to avoid direct competition and to meet diverse and specialized needs of audiences or markets (Carroll and Swaminathan, 2000). While the new institutionalists focused on explaining organizational similarity (i.e., why are organizations so similar?), ecologists focused on variation (i.e., why are organizations so different?). The organizational ecology approach has been echoed by social computing scholars who note that communities can range immensely in their topics, scales, goals, and norms (Chandrasekharan et al., 2018; Weld et al., 2021; TeBlunthuis and Hill, 2022; Zhu et al., 2014, 2013; Hwang and Foote, 2021). Drawing explicitly from organizational ecology, TeBlunthuis et al. (2022) observes how normative differences between communities with similar topics enable communities to be mutualistic rather than competitive with one another. Related work has mapped out “ecologies” of the rules and values that would underpin such normative distinctions (Chandrasekharan et al., 2018; Weld et al., 2021; Fiesler et al., 2018; Nicholson et al., 2023; Hecht and Gergle, 2010). Although researchers have discussed deliberative processes and bureaucratization around rules within specific communities (Black et al., 2008; Butler et al., 2008), we do not know of work that seeks to explain how ecologies of community rules form in online contexts.
@@
An important distinction between the new institutionalism and ecology traditions is in the way they approach change. For new institutionalists, organizational change is a central object of study. For example, a firm may change its rules to comply with a new law or regulation. Similarly, a nonprofit might change its rules to qualify for funding or to conform to published best practices. A key implication from the new institutionalist perspective is that organizations will face similar pressures that will make them structurally similar over time. Conversely, organizational ecologists rely heavily on the concept of structural inertia, which assumes that organizational change is difficult or even impossible (Hannan and Freeman, 1977; Young, 1988). If organizations appear similar, it is because they face similar selection pressures which have caused dissimilar organizations to fail, or because there are positive returns to legitimacy that stem from similar organizations in an environment. That said, some organizational ecologists have argued that certain features of environments may lead to more organizational change (Hannan and Freeman, 1984).
@@
Although the new institutionalist and organizational ecology perspectives are quite different in their assumptions and predictions, both perspectives share a key sociological insight that is largely absent from most scholarship on governance in social computing: the origins of organizational forms—including rules—lie not only in function but also in the relationships between the organizations and their environments. According to both perspectives, we cannot fully understand communities by looking within communities or by focusing on the problems that the rules purport to address. In both perspectives, rules are deeply social and communicative.